Calculate the H3O+ concentration for each pH.

pH=11[H3O+]= 

pH=2[H3O+]= 

pH=5[H3O+]=


1
Expert's answer
2022-04-21T01:07:15-0400

pH = - log10[H3O+] => [H3O+] = 10-pH

pH=11 [H3O+] = 10-11M

pH=2 [H3O+] = 10-2M

pH=5 [H3O+] =10-5M